Circular No. 9116

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

CBAT Director: Daniel W. E. Green; Room 209; Dept. of Earth and Planetary Sciences; Harvard University; 20 Oxford St.; Cambridge, MA 02138, U.S.A. cbatiau@eps.harvard.edu ISSN 0081-0304

                URL http://www.cfa.harvard.edu/iau/cbat.html

COMET P/2010 C1 (SCOTTI) J. V. Scotti, Lunar and Planetary Laboratory, University of Arizona,

reports his discovery of a comet (discovery observation tabulated below) with a nearly stellar nuclear condensation and a 23′′ tail in p.a. 283^{o} on CCD-mosaic images obtained with the Spacewatch 0.9-m f/3 reflector at Kitt Peak.

2010 UT α_{2000} δ_{2000} Mag. Observer
Feb. 9.21136 10^{h}29^{m}06.^{s}76 +11^{o}36′19′′.4 20.2 Scotti

The available astrometry (including single-night observations from Mount Lemmon on 2008 Dec. 30, 2009 Feb. 1, and Feb. 28; and Spacewatch observations from 2009 Jan. 2, 2010 Jan. 12, and Feb. 5), the following elliptical orbital elements, residuals, and an ephemeris appear on MPEC 2010-C26.

Epoch = 2009 Nov. 25.0 TT

T = 2009 Dec. 1.2703 TT  ω = 3.^{o}6353
e = 0.258847  Ω = 142.0341 }2000.0
q = 5.235048 AU  i = 9.1426

a = 7.063383 AU n^{o} = 0.0525031 P = 18.77 years

COMETS C/2009 Q8, C/2009 V1, C/2009 V2 (SOHO) Further to IAUC 9115, additional Kreutz sungrazers have been found

on SOHO website images. K. Battams notes that C/2009 Q8 was small and stellar in appearance (peak mag ∼ 6.5–7). Comets C/2009 V1 and C/2009 V2 were very faint (peak mag ∼ 8–8.5) and slightly diffuse. A. Watson also found images for the latter two comets in STEREO/SECCHI HI-1 website data.

Comet 2009 UT α_{2000} δ_{2000} Inst. F MPEC
C/2009 Q8 Aug. 31.138 10^{h}20.^{m}4 + 7^{o}49′ C3 MU 2009-W10
C/2009 V1 Nov. 5.054 14 32.6 −17 21 C3 MU 2010-B27
C/2009 V2  5.571 14 32.7 −17 53 C3 MU 2010-B27

COMET 29P/SCHWASSMANN-WACHMANN E. Reina, J. Castellano, J. A. Henríquez, and J. M. Aymami report

that F. García (Asturias, Spain) found the comet to be at red mag 11.7 on Feb. 2.999 UT (cf. IAUC 9114). Additional visual total-magnitude estimates: Feb. 6.54, 10.8 (M. Mattiazzo, Castlemaine, Victoria, Australia, 20-cm reflector); 8.10, 10.8 (M. Goiato, Aracatuba, Brazil, 22-cm reflector).

2010 February 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2010 CBAT  Daniel W. E. Green